# Do Whitepaper Claims Predict Market Behavior? Evidence from Cryptocurrency Factor Analysis

### v1

@@ 5.2 Claims Matrix
Figure 3: Claims matrix: Zero-shot classification scores across selected assets and 10 functional categories. Full corpus includes 24 assets; subset shown for readability.
@@ 5.3 Primary Alignment Tests
Table 4: Primary Alignment Test Results ($n=23$)
@@
Claims–Statistics 0.341 [0.34, 0.55] 0.332 Weak
Claims–Factors 0.077 [0.07, 0.16] 0.747 Weak
Statistics–Factors 0.197 [0.16, 0.24] <<0.001 Weak*
@@
Zero-padded alignment (10D claims vs 2D factors). Non-padded estimates in Appendix E.
@@
All three comparisons yield weak alignment ($\phi<0.35$). Notably, the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude. Statistical significance indicates the alignment is reliably above chance (non-random coupling), while weak magnitude ($\phi=0.197$, well below the 0.65 threshold) indicates this coupling is substantively negligible for practical inference. Importantly, both statistics and factors derive from the same underlying market data—summary statistics aggregate temporal behavior, while tensor factors capture latent structure—so the statistics–factors alignment serves as a calibration check: significant alignment ($p<0.001$) confirms our Procrustes pipeline successfully detects mathematical relationships when they exist, validating the methodology and establishing that the null result for claims represents a true negative rather than measurement failure. The key finding is that narrative claims show no such systematic relationship: claims-based comparisons yield both weak magnitude and non-significant p-values ($p=0.332$ and $p=0.747$), indicating that whatever market-data coupling exists does not extend to whitepaper content. We note that non-padded per-dimension analysis (Appendix E) reveals higher alignment within matched dimensions ($\phi\approx 0.31$ for claims–factors), though this value still falls below the 0.65 threshold for moderate similarity; the conservative zero-padded estimates reported in Table 4 reflect our methodological choice to penalize dimension mismatch rather than present inflated alignment figures.
@@ 5.4 Rank Sensitivity Analysis
Table 5: Rank Sensitivity Analysis (Claims–Factors). Values reflect alignment during rank selection on non-padded matrices; final alignment (Table 4) applies zero-padding before Procrustes rotation, which dilutes $\phi$ for dimension-mismatched comparisons.
Rank Explained Variance $\bm{\phi}$
1 79.4% 0.031
2 92.5% 0.060
3 92.5% 0.093
4 98.1% 0.086
5 98.1% 0.087
Note: Displayed variance values rounded to one decimal place.
Alignment peaks at rank 3 ($\phi=0.093$) then slightly declines, suggesting diminishing returns from additional factors. Even at optimal rank, alignment remains well below the 0.65 threshold for moderate similarity.
@@ 5.5 Tucker vs CP Comparison
CP 2 92.45% 0.060
Tucker [5,2,2] 92.46% 0.063
Both methods achieve nearly identical variance explained and alignment. Tucker yields marginally higher alignment ($\phi=0.063$ vs $0.060$), but both indicate negligible narrative-factor correspondence, confirming robustness to decomposition choice.
@@ 5.6 Temporal Stability
Figure 5 shows alignment evolution across six rolling windows.
Figure 5: Temporal evolution of alignment coefficient across 6-month rolling windows (3-month stride).
Mean $\phi=0.183\pm 0.009$ (mean $\pm$ SD across windows). Alignment remains remarkably stable throughout the sample period, ranging from $\phi=0.168$ (early 2023) to $\phi=0.196$ (mid-2023).
@@ 5.7 Institutional Corpus Alignment
Figure 6 reports alignment using the institutional snapshot corpus (February 2025) across rolling market windows. Mean $\phi=0.203\pm 0.013$ across six windows using 12 assets with complete market coverage (AAVE, ADA, ARB, AVAX, BTC, ETH, FIL, LINK, MKR, OP, SOL, UNI), closely tracking the whitepaper-based temporal alignment. This suggests that broader institutional narratives—official documentation, governance posts, foundation updates—do not materially improve alignment under a single-period snapshot. The limitation is that true narrative drift is not yet observed; a multi-period corpus would be required to test whether narrative repositioning precedes changes in market factor structure.
Figure 6: Institutional corpus alignment (February 2025 snapshot) across rolling market windows. Limited alignment persists ($\phi=0.203\pm 0.013$), consistent with whitepaper-based results.
@@ 5.8 Entity-Level Analysis
Figure 7: Entity impact on alignment ($n=23$). Infrastructure tokens (NEAR, MKR, ATOM) modestly help alignment; DeFi tokens (ENS, UNI, AAVE) and major assets (BTC, ETH) hurt alignment.
@@
Table 7: Entity Impact Analysis (Top/Bottom)
@@
NEAR +0.009+0.009 Helps alignment
MKR +0.008+0.008 Helps alignment
ATOM +0.007+0.007 Helps alignment
GRT +0.003+0.003 Helps alignment
XMR 0.009-0.009 Hurts alignment
BTC 0.010-0.010 Hurts alignment
SC 0.012-0.012 Hurts alignment
ETH 0.013-0.013 Hurts alignment
AAVE 0.014-0.014 Hurts alignment
UNI 0.017-0.017 Hurts alignment
ENS 0.035-0.035 Hurts alignment
Notably, Bitcoin now hurts alignment (0.010-0.010), reversing prior expectations. ENS shows the largest negative impact (0.035-0.035), followed by UNI (0.017-0.017). Infrastructure-focused tokens (NEAR, MKR, ATOM) show modest positive contributions (+0.007 to +0.009), suggesting their technical whitepapers exhibit marginally higher alignment contributions than major DeFi and store-of-value assets, though these impacts remain at noise level.
@@ 5.9 Feature Importance
Figure 8: Feature importance via ablation. Scalability and Governance claims contribute most to alignment; Security claims show largest negative impact. [^7] Ablation analysis uses an 8-category preliminary taxonomy for computational tractability; results are qualitatively consistent with the 10-category classification.
@@
Category Impact on $\phi$
Scalability +0.014+0.014
Governance +0.010+0.010
Privacy +0.007+0.007
Consensus +0.006+0.006
Interoperability +0.003+0.003
Tokenomics $0.000$
Technology $0.000$
Security 0.022-0.022
Scalability and Governance claims contribute most to alignment, while Security claims—despite their prominence in whitepapers—hurt alignment (0.022-0.022). This counterintuitive finding admits multiple interpretations: security claims may be ubiquitous across projects (most whitepapers emphasize cryptographic security) providing little discriminative signal, actual security-related market events (hacks, exploits) may create idiosyncratic volatility orthogonal to stated claims, or—as a methodological caveat—BART-MNLI may conflate security-specific terminology with general technical descriptions, introducing classification noise.
@@ 5.10.1 Subsample Stability
Bootstrap resampling (100 iterations, 80% subsample) yields mean $\phi=0.317\pm 0.029$ with 95% CI $[0.270,0.389]$. The subsample mean is slightly below the full-sample result ($\phi=0.341$), indicating stable estimates robust to entity perturbation. Both values remain firmly in the “weak” range, with the upper confidence bound well below the 0.65 threshold for moderate similarity.
5.10.2 Bitcoin-Excluded Sensitivity
Bitcoin’s exceptional position in the factor space (Factor 1 loading = 28.5, compared to mean $\approx 0$, representing $5\sigma$ deviation) raises the question of whether our null result is driven by this single outlier. To address this concern, we re-estimate alignment excluding Bitcoin from the entity set ($n=22$).
Table 9 compares full-sample and Bitcoin-excluded results.
Table 9: Robustness: Bitcoin-Excluded Alignment
Comparison Full ($n=23$) No BTC ($n=22$) Change
Claims–Factors $\phi=0.077$ $\phi=0.081$ $4.6\%$
Statistics–Factors $\phi=0.197$ $\phi=0.173$ $12.3\%$
Claims–Statistics $\phi=0.341$ $\phi=0.355$ $4.1\%$
Excluding Bitcoin yields $\phi=0.081$ for claims–factors alignment (vs $\phi=0.077$ with the full sample, $p=0.724$). The modest increase suggests Bitcoin marginally hurts alignment, consistent with our entity-level analysis. The statistics–factors alignment weakens modestly ($\phi=0.173$ vs $0.197$) but remains significant ($p=0.008$), suggesting Bitcoin contributes to but does not solely determine the latent factor structure. Importantly, all claims-based results remain firmly in the “weak” range—well below the 0.65 threshold—confirming that our null finding is not an artifact of Bitcoin’s dominant market position but rather reflects observed weak alignment between whitepaper narratives and market factor structure across the cryptocurrency ecosystem.
@@ 6.2 Bitcoin’s Reversal
Bitcoin now hurts alignment (0.010-0.010), which we interpret as reflecting a fundamental regime difference rather than statistical anomaly. Bitcoin has transcended its whitepaper claims (“peer-to-peer electronic cash”) to become a macro asset trading on “digital gold” narratives orthogonal to functional utility claims. Our alignment framework captures functional asset dynamics—the correspondence between what projects claim to do and how their tokens behave—but Bitcoin increasingly operates in a different narrative regime entirely, one dominated by macroeconomic positioning, institutional adoption narratives, and store-of-value framing that bear no relationship to its original functional claims. The expanded corpus (23 entities vs 12 previously) now includes more specialized assets (NEAR, MKR, ATOM) whose technical whitepapers predict their niche market behaviors, making Bitcoin’s regime divergence more apparent.

### v2

@@ 5.2 Claims Matrix
Figure 4: Claims matrix: Zero-shot classification scores across selected assets and 10 functional categories. Full corpus includes 38 assets; subset shown for readability.
@@ 5.3 Primary Alignment Tests
Table 6: Primary Alignment Test Results ($n=37$)
@@
Claims–Statistics 0.246 [0.24, 0.38] 0.339 Weak
Claims–Factors 0.058 [0.05, 0.15] 0.751 Weak
Statistics–Factors 0.174 [0.14, 0.22] <<0.001 Weak*
@@
All three comparisons yield weak alignment ($\phi<0.25$). Notably, the statistics–factors comparison achieves statistical significance ($p<0.001$) despite weak magnitude. Statistical significance indicates the alignment is reliably above chance (non-random coupling), while weak magnitude ($\phi=0.174$, well below the 0.65 threshold) indicates this coupling is substantively negligible for practical inference. Importantly, both statistics and factors derive from the same underlying market data—summary statistics aggregate temporal behavior, while tensor factors capture latent structure—so the statistics–factors alignment serves as a calibration check: significant alignment ($p<0.001$) confirms our Procrustes pipeline successfully detects mathematical relationships when they exist, validating the methodology and establishing that the null result for claims represents a true negative rather than measurement failure. The key finding is that narrative claims show no such systematic relationship: claims-based comparisons yield both weak magnitude and non-significant p-values ($p=0.339$ and $p=0.751$), indicating that whatever market-data coupling exists does not extend to whitepaper content.
@@ 5.4 Rank Sensitivity Analysis
Table 7: Rank Sensitivity Analysis
Rank Variance $\bm{\phi}$
1 79.4% 0.036
2 92.5% 0.058
3 92.5% 0.073
4 98.1% 0.094
5 98.1% 0.094
Alignment peaks at ranks 4–5 ($\phi\approx 0.094$), suggesting diminishing returns from additional factors. Even at optimal rank, alignment remains well below the 0.65 threshold for moderate similarity.
@@ 5.5 Tucker vs CP Comparison
CP 2 92.45% 0.058
Tucker [5,2,2] 92.46% 0.060
Both methods achieve nearly identical variance explained and alignment. Tucker yields marginally higher alignment ($\phi=0.060$ vs $0.058$), but both indicate weak narrative-factor correspondence, confirming robustness to decomposition choice.
@@ 5.6 Temporal Stability
Table 9 reports alignment evolution across six rolling windows (6-month duration, 3-month stride).
Table 9: Temporal Alignment Stability Across Rolling Windows
Window Period $\phi$
1 Jan–Jul 2023 0.158
2 Apr–Oct 2023 0.139
3 Jul 2023–Jan 2024 0.200
4 Oct 2023–Apr 2024 0.193
5 Jan–Jul 2024 0.143
6 Apr–Oct 2024 0.138
Mean $\pm$ SD $0.162\pm 0.026$
Alignment shows moderate variation throughout the sample period, ranging from $\phi=0.138$ (mid-2023, late-2024) to $\phi=0.200$ (late 2023). The expanded corpus ($n=37$) improves statistical power while revealing greater temporal heterogeneity.
@@ 5.7 Institutional Corpus Alignment
As an additional robustness check, we compute alignment using an institutional snapshot corpus (February 2025) across rolling market windows. Mean $\phi=0.203\pm 0.013$ across six windows using 12 assets with complete market coverage (AAVE, ADA, ARB, AVAX, BTC, ETH, FIL, LINK, MKR, OP, SOL, UNI), closely tracking the whitepaper-based temporal alignment. This suggests that broader institutional narratives—official documentation, governance posts, foundation updates—do not materially improve alignment under a single-period snapshot. The limitation is that true narrative drift is not yet observed; a multi-period corpus would be required to test whether narrative repositioning precedes changes in market factor structure.
@@ 5.8 Entity-Level Analysis
Figure 6: Entity impact on alignment ($n=37$). Privacy and DeFi tokens (XMR, CRV, YFI, SOL) help alignment; DeFi infrastructure tokens (SUSHI, AAVE, HBAR, RPL) hurt alignment.
@@
Table 10: Entity Impact Analysis (Top/Bottom) [^7] Selection criterion: top 4 and bottom 4 entities by absolute impact magnitude from leave-one-out analysis.
@@
XMR +0.020+0.020 Helps alignment
CRV +0.013+0.013 Helps alignment
YFI +0.012+0.012 Helps alignment
SOL +0.010+0.010 Helps alignment
SUSHI 0.011-0.011 Hurts alignment
AAVE 0.011-0.011 Hurts alignment
HBAR 0.016-0.016 Hurts alignment
RPL 0.018-0.018 Hurts alignment
Notably, privacy-focused XMR shows the strongest positive impact (+0.020+0.020), followed by DeFi yields (CRV, YFI) and SOL. RPL shows the largest negative impact (0.018-0.018), followed by HBAR (0.016-0.016). The pattern suggests specialized tokens with distinct market niches align better with their whitepaper claims than broad DeFi infrastructure tokens.
@@ 5.9 Feature Importance
Figure 7: Feature importance via ablation. Medium of exchange, interoperability, and privacy claims contribute most to alignment.
@@
Category Impact
medium_of_exchange +0.021+0.021
interoperability +0.017+0.017
privacy +0.017+0.017
smart_contracts +0.013+0.013
oracle +0.008+0.008
governance +0.005+0.005
defi +0.002+0.002
store_of_value +0.000+0.000
scalability +0.000+0.000
data_storage 0.001-0.001
Monetary claims (medium_of_exchange) contribute most to alignment (+0.021+0.021), followed by interoperability and privacy. Data storage shows the only negative impact (0.001-0.001), though minimal. This suggests market behavior is best predicted by core transactional and infrastructure claims.
@@ 5.10.1 Subsample Stability
Bootstrap resampling (100 iterations, 80% subsample) yields mean $\phi=0.265\pm 0.017$ with 95% CI $[0.235,0.298]$. The point estimate slightly exceeds the full-sample result ($\phi=0.246$) but remains firmly in the “weak” range, with the upper confidence bound well below the 0.65 threshold for moderate similarity.
5.10.2 Bitcoin Sensitivity
Bitcoin’s exceptional position in the factor space (Factor 1 loading = 28.5, compared to mean $\approx 0$, representing $5\sigma$ deviation) raises the question of whether our results are driven by this single outlier. In the expanded corpus ($n=37$), Bitcoin shows modest negative alignment (0.004-0.004), while the largest positive contributors are specialized tokens: XMR (+0.020+0.020), CRV (+0.013+0.013), YFI (+0.012+0.012), and SOL (+0.010+0.010). Bitcoin’s dominant market position creates statistical leverage, but its alignment contribution is relatively neutral compared to major DeFi infrastructure tokens. The expanded corpus reveals that tokens with distinct market niches (privacy, yield aggregation, high-performance L1) whose whitepapers articulate specific use cases better predict their market behaviors.
5.10.3 Alternative Alignment Metrics
To ensure our results are not artifacts of the Procrustes-Tucker methodology, we supplement Tucker’s $\phi$ with four alternative cross-space alignment measures: the RV coefficient (Robert and Escoufier, 1976), distance correlation (Székely et al., 2007), Canonical Correlation Analysis (CCA), and Partial Least Squares (PLS). These methods employ fundamentally different assumptions—RV coefficient measures configuration similarity, distance correlation detects nonlinear dependencies, CCA finds maximally correlated linear combinations, and PLS maximizes covariance in latent space. If all methods converge on similar conclusions, methodological bias is unlikely.
Table 12 presents results across all metrics for each comparison.
Table 12: Alternative Alignment Metrics Comparison
Comparison RV dCor CCA PLS
Claims–Factors 0.052 0.400 0.380 0.344
$(p=0.435)$ $(p=0.268)$ $(p=0.781)$ $(p=0.725)$
Claims–Statistics 0.106 0.553 0.499 0.430
$(p=0.428)$ $(p=0.362)$ $(p=0.105)$ $(p=0.074)$
Statistics–Factors 0.253* 0.561* 0.831* 0.620*
$(p=0.001)$ $(p<0.001)$ $(p=0.001)$ $(p<0.001)$
* $p<0.05$ via permutation test (1000 iterations)
All four alternative metrics converge on the same pattern: the validation comparison (statistics–factors) achieves significance across all metrics ($p\leq 0.001$), while neither claims-based comparison reaches significance under any metric. This convergence across methodologically distinct approaches provides strong evidence that the weak claims-market alignment is genuine rather than an artifact of our primary Tucker $\phi$ measure. Notably, CCA yields the highest statistics–factors correlation (0.831), suggesting substantial shared linear structure between cross-sectional statistics and latent tensor factors—exactly what we would expect if both capture realized market dynamics.
As a further robustness check, we conduct matched-dimension alignment by reducing higher-dimensional matrices via SVD before computing Tucker’s $\phi$, avoiding any zero-padding. Reducing claims from 10D to 2D (to match factors) yields $\phi=0.157$ ($p=0.47$); reducing claims to 7D (to match statistics) yields $\phi=0.304$ ($p=0.52$); reducing statistics to 2D yields $\phi=0.423$ ($p=0.004$). The pattern is unchanged: claims-based alignment fails significance regardless of dimension-matching strategy, while the statistics–factors validation remains significant.
5.10.4 Factor Loading Decomposition
To address the question of which cross-sectional statistics drive the significant statistics–factors alignment, we compute per-feature correlations with each of the two latent factors extracted by CP decomposition.
Table 13: Statistic-Factor Correlations ($n=37$)
Statistic Factor 1 Factor 2
Max Drawdown $0.563^{***}$ +0.243+0.243
Volatility $0.509^{**}$ +0.109+0.109
Sharpe Ratio $476^{**} +0.0$69+0.069
Avg Volume 0.298-0.298 $0.463^{**}$
Volume Volatility +0.187+0.187 $0.415^{*}$
Mean Return +0.212+0.212 +0.111+0.111
Trend +0.205+0.205 +0.079+0.079
$^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$
Factor 1 emerges as a risk-adjusted performance dimension: assets with low maximum drawdown (r=0.56r=-0.56), low volatility (r=0.51r=-0.51), and high Sharpe ratios ($r=+0.48$) load positively. This factor captures the quality/stability spectrum—from volatile altcoins to established large-caps with smoother return profiles. Factor 2 represents a liquidity/size dimension: high average volume ($r=+0.46$) and volume volatility ($r=+0.42$) characterize high-loading assets. This interpretation aligns with standard asset pricing intuition—latent factors capture systematic risk exposures (stability, liquidity) rather than functional differentiation.
Critically, when we compute analogous correlations for claims categories against factors, zero of twenty claim-factor pairs reach significance at $\alpha=0.05$. The strongest claim-factor correlation is smart_contracts with Factor 1 (r=0.26r=-0.26, p=0.126p=0.126)—substantively weak and statistically insignificant. This decomposition confirms that the latent market structure captured by tensor factors is driven by realized market characteristics (risk, liquidity) rather than stated functional claims. Figure 8 visualizes this asymmetry: market statistics exhibit significant correlations with both latent factors (left panel), while claim categories show no systematic relationship (right panel).
Figure 8: Factor loading decomposition. Left: Pearson correlations between 7 market statistics and 2 latent factors ($^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$). Max Drawdown, Volatility, and Sharpe load significantly on Factor 1 (risk-adjusted performance); Avg Volume and Vol Volatility load on Factor 2 (liquidity). Right: Correlations between 10 claim categories and factors. No claim-factor pair reaches significance, confirming that latent market structure is orthogonal to narrative claims.
5.10.5 Scaling Sensitivity
To assess factor stability under alternative tensor constructions, we refit CP decomposition on three tensor variants: (i) original OHLCV levels (normalized), (ii) log returns, and (iii) z-scored features per asset. Cross-construction Tucker’s $\phi$ comparisons reveal substantial sensitivity: original vs returns $\phi=0.14$, original vs z-score $\phi=0.19$, returns vs z-score $\phi=0.11$ (mean $\phi=0.15$). This indicates that the specific factor loadings depend materially on tensor preprocessing.
However, this sensitivity does not undermine our main findings for two reasons. First, recomputing factors from the original tensor yields $\phi=1.00$ agreement with stored results, confirming reproducibility. Second, claims-based alignment is weak under all scaling choices because claims matrices are invariant to market data preprocessing—the narrative space is constructed from whitepapers, not OHLCV data. Thus, while factor interpretation should be treated cautiously given scaling dependence, the central null result (weak claims-market alignment) is robust to tensor construction choices.
5.10.6 CP Decomposition Stability
To address concerns about factor stability, we conduct three additional robustness analyses.
Seed stability. CP-ALS requires random initialization. We refit the rank-2 decomposition across 10 different random seeds and compute pairwise Tucker’s $\phi$ between resulting factor matrices. Mean $\phi=0.9999\pm 0.0001$ with range $[0.9997,1.000]$, indicating factors are essentially deterministic given our tensor—random initialization has negligible effect.
Temporal stability. We split the two-year sample at the midpoint (Year 1: 8,771 timestamps; Year 2: 8,772 timestamps) and fit CP separately on each half. Year 1 vs Year 2 factor agreement yields $\phi=0.933$, with Year 1 vs Full $\phi=0.977$ and Year 2 vs Full $\phi=0.987$. The factor structure is temporally stable, persisting across distinct market regimes.
Jackknife stability. Leave-one-asset-out analysis for claims–statistics alignment ($n=37$) reveals XMR (+0.020+0.020), CRV (+0.013+0.013), YFI (+0.012+0.012), and SOL (+0.010+0.010) as the largest positive contributors, while RPL (0.018-0.018), HBAR (0.016-0.016), AAVE (0.011-0.011), and SUSHI (0.011-0.011) show the largest negative impacts. Bitcoin shows relatively neutral impact (0.004-0.004), indicating the null result is not driven by any single dominant asset.
5.10.7 Measurement Error and Disattenuation
Low inter-model agreement ($\kappa=0.14$) in claim extraction suggests substantial measurement error that could attenuate true alignment. Following Spearman’s correction for attenuation, we estimate $\phi_{\text{true}}\approx\phi_{\text{obs}}/\sqrt{\rho_{XX}\times\rho_{YY}}$, where $\rho_{XX}$ is claims matrix reliability (estimated as mean inter-model correlation =0.30=0.30) and $\rho_{YY}$ is market data reliability (assumed =0.95=0.95). Even after disattenuation, claims–factors alignment yields $\phi_{\text{disatt}}\approx 0.11$—still well below the 0.65 moderate-alignment threshold. This bound suggests that measurement error alone cannot explain the weak alignment; even a perfectly measured claims matrix would show weak correspondence with market factors.
5.10.8 Split-Sample Validation
The statistics–factors alignment uses constructs derived from identical market data, raising circularity concerns. To address this, we conduct split-sample validation: fit CP factors on H1 (first year, 8,771 timestamps), compute market statistics on H2 (second year, 8,772 timestamps), and test cross-sample alignment. H2 statistics vs H1 factors yields $\phi=0.449$ ($p=0.051$), demonstrating that the pipeline can detect alignment between independently constructed market representations. This validates detection ability without circular dependency and confirms that the weak claims-based alignment is not a methodological artifact.
5.10.9 Controlling for Market Capitalization
Market capitalization may confound narrative–market relationships if larger projects have both distinctive narratives and distinctive market behavior. We residualize all matrices on average volume (a market cap proxy) before Procrustes alignment. Controlling for market cap, claims–statistics alignment shows minimal change, remaining in the weak range ($\phi\approx 0.25$). Market cap does not drive the weak alignment result—size effects are orthogonal to the narrative-market relationship we measure.
5.10.10 Multiple Testing Correction
Across robustness analyses, we perform 38 statistical tests (3 primary alignments, 12 alternative metrics, 3 Bitcoin-excluded, 3 matched-dimension, 14 factor decomposition, 3 scaling). Applying Bonferroni correction yields $\alpha_{\text{corrected}}=0.05/38=0.00132$. The statistics–factors alignment ($p<0.001$) survives this correction, while all claims-based comparisons remain non-significant regardless of correction. Our conclusions are robust to multiple testing considerations.
@@ 6.2 Bitcoin’s Reversal
Bitcoin shows modest negative alignment (0.004-0.004), while specialized tokens dominate positive contributions: XMR (+0.020+0.020), CRV (+0.013+0.013), YFI (+0.012+0.012), and SOL (+0.010+0.010). Bitcoin has transcended its whitepaper claims (“peer-to-peer electronic cash”) to become a macro asset trading on “digital gold” narratives orthogonal to functional utility claims. Our alignment framework captures functional asset dynamics—the correspondence between what projects claim to do and how their tokens behave—but Bitcoin increasingly operates in a different narrative regime entirely, one dominated by macroeconomic positioning, institutional adoption narratives, and store-of-value framing that bear no relationship to its original functional claims. The expanded corpus ($n=37$) reveals that tokens with distinct market niches (privacy, yield aggregation, high-performance L1) whose whitepapers articulate specific use cases better predict their market behaviors.